OpenStudy (anonymous):

I'm having such a hard time understand some questions i read my lesson twice looked up youtube video's and my teacher is unavailable so i need help. Solve the literal equation for the unknown variable. So rewrite the formula as w =

OpenStudy (anonymous):

\[p = 2l + 2w\]

OpenStudy (anonymous):

@DrummerGirl3

OpenStudy (anonymous):

@jabez177

OpenStudy (drummergirl3):

u need to solve for L p=2L+2w Or 2L + 2W = P Subtract 2W from both sides and you have: 2L = P - 2W Divide both sides by 2 L = (P-2W)/2 or L = P/2 - W

OpenStudy (drummergirl3):

Got it?

OpenStudy (anonymous):

im actually trying to solve for w.

OpenStudy (drummergirl3):

ohh lmao